somebody said the solution for:squareroot of negative w, then subracting 6=w is negative 4 and negative 9.Then he said we take negative 9 because the squareroot of 9 (after substituding w with -9) is -3.I am confussed.please explain this to me.
Answer by RAY100(1637)

sqrt(-w) -6 =w
sqrt(-w) = w+6
SQUARE BOTH SIDES
(-w) = (w+6)^2 = w^2 +12w +36
0 = w^2 +13w +36
FACTOR
0 = ( w+9)(w+4)
(w+9) =0,,,w=(-9),,,,,,,(w+4)=0,,,,,w=(-4),,,,so far ok to advise
.
Checking
w=-4,,,,,sqrt -(-4) -6 = (-4)
sqrt(+4) -6 = -4
+2 -6 = -4,,,,,,ok,,,,,,,,,,,,,,,,,,pls note,,,,,,opposite advise
.
w=(-9),,,,,,sqrt -(-9) -6 = (-9)
sqrt (+9) -6 = (-9)
3 - 6 = -9
-3 = -9,,,,,NO,,,,,,therefore not answer,,,,,note,,,opposite advise